QUADRATIC
2x^2+3x=9

Answer by richwmiller(17219)


2x^2+3x=9
2x^2+3x-9=0
You want factors of -18 (2*-9) which add up to 3
6 and -3
2x^2+6x-3x-9
2x(x+3)-3*(x+3)
(2x-3)*(x+3)=0
2x-3=0
2x=3
x=3/2
x+3=0
x=-3